What is the momentum (in MeV/c) and kinetic energy of a proton with a total energy of 1.00 GeV? Repeat for an electron with the same energy.
Answer #1

What is the momentum (in MeV/c) and kinetic energy of a proton with a total energy of 1.00 GeV?

Solution :-

the energy of aproton E =√[p2c2 +(mc2)2]

For proton mc2 =(1.67*10-27kg)(3*108m/s)2

                        = 15.03*10-11J

                         =(15.03*10-11J)/(1.6*10-10GeV/J)

                         =  0.94GeV

Then pc = √[(1GeV)2 -(0.94GeV)2]

            = 0.3412GeV

or momentum of proton p = 0.3412GeV/c

Kinetic energy K = E - mc2

                          = 1GeV - 0.94GeV

                          = 0.06GeV

Do it for electron by putting m = 9.1*10-31kg